OpenStudy (anonymous):

6(x - 1) = 4(x - 2) is it -1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6x-6=4x-8 6x-4x=-8+6 2x= -2 x=-2/2 x=-1

OpenStudy (anonymous):

yeah